Pentane is burnt with 96000cm^3 of oxygen gas at room temperature and pressure.

i. Write the balanced chemical equation for the reaction.

ii. Calculate the volume of carbon dioxide in cubic centimetres (cm^3) produced in the reaction at room temperature and pressure.

iii. Calculate the mass of Zinc needed to react with 50cm^3 0.8M of Hydrochloric acid.


1
Expert's answer
2021-10-08T02:09:46-0400

C5H12+8O2→6 H2O+5 CO2

Molar mass of Oxygen = 31.998


96000/31.998


= 3000.19m


Molar mass of CO2 = 44.01 g/mol


3000.19× 44.01× 5/8


= 82523.98cm3


iii.Zn + HCl = H2 + ZnCl2

Molar mass of HCl = 36.458 g/mol

0.8× 36.458= 29.1664g

50/29.1664= 1.714mol

Mole ratio = 1:1

Molar mass of Zn = 65.38g/mol

1.714×65.38= 112.08g